Dividing and Multiplying Integers 1) 6x-4 2)6/-1 3) 3x-4 4) 65/-13 5)5X-4 6) -168/-12 7) -5X6
1
Expert's answer
2013-07-03T08:17:10-0400
1) 6x-4=-24 & & 2)6/-1=-6 & & 3) 3x-4=-12 & & 4) 65/-13=-5 & & 5)5X-4=-20 & 6) -168/-12=14 & & 7) -5X6=-30&